# L-PR: Exploiting LiDAR Fiducial Marker for Unordered Low Overlap Multiview Point Cloud Registration

### v2

@@ Abstract
Point cloud registration is a prerequisite for many applications in computer vision and robotics. Most existing methods focus on pairwise registration of two point clouds with high overlap. Although there have been some methods for low overlap cases, they struggle in degraded scenarios. This paper introduces a novel framework dubbed L-PR, designed to register unordered low overlap multiview point clouds leveraging LiDAR fiducial markers. We refer to them as LiDAR fiducial markers, but they are the same as the popular AprilTag and ArUco markers—thin sheets of paper that do not affect the 3D geometry of the environment. We first propose an improved adaptive threshold marker detection method to provide robust detection results when the viewpoints among point clouds change dramatically. Then, we formulate the unordered multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop a framework consisting of two levels of graphs to address it. The first-level graph, constructed as a weighted graph, is designed to efficiently and optimally infer initial values of scan poses from the unordered set. The second-level graph is constructed as a factor graph. By globally optimizing the variables on the graph, including scan poses, marker poses, and marker corner positions, we tackle the MAP problem. We conduct both qualitative and quantitative experiments to demonstrate that the proposed method surpasses previous state-of-the-art (SOTA) methods and to showcase that L-PR can serve as a low-cost and efficient tool for 3D asset collection and training data collection. In particular, we collect a new dataset named Livox-3DMatch using L-PR and incorporate it into the training of the SOTA learning-based method, SGHR, which brings evident improvements for SGHR on various benchmarks. We release the open-source code implementation and datasets at https://github.com/yorklyb/L-PR.
@@ I Introduction
Point cloud registration is a fundamental need in computer vision [1, 2] and robotics [3, 4, 5]. Most existing point registration methods [3, 6] focus on pairwise registration of two point clouds with high overlap, which are not robust in practical applications [2]. Some recent learning-based research proposes multiview registration methods [1, 7] and studies low overlap cases [2, 1]. Despite their promising performance on benchmarks, the generalization of learning-based methods to unseen scenarios (i.e., out-of-distribution cases of training data) remains problematic. Moreover, both geometry-based methods [3, 4, 5, 8] and learning-based methods [1, 2] rely on geometric features in the scenes to align point clouds. Hence, these methods struggle in degraded scenes [9], such as those with repetitive structures and weak textures.
Reviewing the domain of 2D images, it is found that fiducial markers, such as AprilTag [10] and ArUco [11], have been widely used to tackle challenging situations, such as dramatic viewpoint changes (low overlap) between frames and textureless environments [12, 13, 14]. The visual fiducial markers have rich patterns that are robust for detection and matching, thanks to elaborately designed encoding-decoding algorithms [10, 11]. They pose as thin sheets of paper that can be attached to surfaces without disrupting the 3D environment. In contrast, fiducial objects developed for LiDARs, such as calibration boards [15, 16, 17], lack encoding-decoding pattern design and often take the form of large boards mounted on tripods, thus affecting the 3D environment. LiDARTag [18] is the first work to develop a fiducial marker system akin to AprilTag [10] for LiDARs. While its patterns are compatible with AprilTag, the marker itself is still an additional object placed on a tripod. Recently, the intensity image-based LiDAR fiducial marker system (IFM) [19] proposes to utilize intensity and range images to detect markers. Thus, IFM is compatible with popular visual fiducial markers [10, 11] and inherits the virtue of not affecting the 3D environment. However, the marker detection of IFM is not robust when the viewpoints change in a wide range [19], which hinders its application to in-the-wild multiview point cloud registration. Although the development of LiDAR fiducial markers [18, 19] has advanced, their applications are limited to calibration [20] and AR [19]. The utilization of LiDAR fiducial markers in multiview point cloud registration remains an open problem.
@@
In this paper, we develop a novel framework, dubbed L-PR (see Fig. 1), to exploit LiDAR fiducial markers for unordered low overlap multiview point cloud registration. L-PR is a new solution for training data collection that only uses a LiDAR sensor. As seen, the LiDAR fiducial markers are thin sheets of board/paper attached to other surfaces without impacting the environment. Specifically, the LiDAR fiducial markers can be printed on paper using standard printers, offering a low-cost, easy-to-use, and reliable alternative compared to previous external sensors.
Figure 1: (a): Overview of L-PR. (1) Given an unordered set of unaligned, low overlap 3D point clouds, we aim to register them into (5) a complete point cloud utilizing LiDAR fiducial markers (thin sheets of paper/board attached to other planes). This is achieved through the proposed (2) adaptive threshold marker detection method and the (3)(4) two-level graphs. The (3) first-level graph handles the unordered set of point clouds and calculates initial values. The (4) second-level graph registers the point clouds by finding the optimal solution to a maximum a-posteriori problem. (b): The proposed method is robust to degraded scenes, such as scenes with repetitive structures.
Given an unordered set of low overlap point clouds, the proposed L-PR efficiently registers them in a global frame. To address the unstable marker detection of IFM [19] in the wild, we first develop an adaptive threshold marker detection method that is robust to viewpoint switching. Then, we formulate the multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop two-level graphs to tackle it. The first-level graph, constructed as a weighted graph, is designed to efficiently process the unordered set of point clouds and calculate the initial values of poses among point clouds. In particular, the weights represent the pose estimation error of each marker observation, and we obtain the optimal solutions of initial values by finding the shortest path from an anchor scan to each non-anchor scan on the graph. Given the initial values, we design the second-level graph, a factor graph, to resolve the MAP problem by globally optimizing the poses of point clouds, markers, and positions of marker corners. We conduct both qualitative and quantitative experiments to demonstrate the superiority of the proposed method over competitors [1, 3, 14, 4] and to showcase the diverse applications of L-PR, including 3D asset collection, training data collection, reconstruction of a degraded scene, localization in a GPS-denied environment, and 3D map merging. We release the code implementation and datasets at https://github.com/yorklyb/L-PR.
@@
  • We design a novel framework, named L-PR, for unordered low overlap multiview point cloud registration leveraging LiDAR fiducial markers. L-PR can serve as an efficient and reliable tool to collect 3D assets and training data for learning-based methods.
  • We collect a new training dataset called Livox-3DMatch using the proposed L-PR. Livox-3DMatch augments the original 3DMatch training data [6] from 14,400 pairs to 17,700 pairs (a 22.91% increase). By training on this augmented dataset, the performance of the SOTA learning-based method SGHR [1] is improved by 2.90% on 3DMatch [6], 4.29% on ETH [21], and 22.72% (translation) / 11.19% (rotation) on ScanNet [22].
@@ II Related Works
LiDAR Fiducial Markers. Given the success of Structure from Motion [12, 14] and visual Simultaneous Localization and Mapping (SLAM) [13, 23] approaches based on visual fiducial markers [10, 11], why do we still need to exploit LiDAR and LiDAR fiducial markers? There are three reasons. (1) Unlike image-based fiducial marker detection, which is sensitive to ambient light, LiDAR fiducial marker detection is robust to unideal illumination conditions [18, 19], such as purely dark or overexposed environments. (2) Unlike image-based fiducial marker detection, which suffers from rotational ambiguity [12, 14, 13] and requires abandoning ambiguous measurements, LiDAR fiducial marker pose estimation is not affected by rotational ambiguity [18, 19] and each marker observation can be utilized. (3) Data captured by LiDARs is irreplaceable by data acquired from cameras due to the unique modality of LiDAR point cloud [24].
@@
Point Cloud Registration. Geometry-based point cloud registration methods, such as variants of iterative closest point methods [4, 5], RANSAC [8], and Teaser++ [3], mainly focus on designing efficient and robust algorithms for describing geometry and extracting geometric features (e.g., points, lines, and surfaces/normals) to find correspondence between point clouds and align them. In learning-based point cloud registration methods like Predator [2] and SGHR [1], neural networks are designed to learn features representing the geometry [25, 26, 27, 7, 6], and then learn to utilize these features for registering the point clouds. While these methods demonstrate remarkable performance on benchmarks [6, 21, 22], they still face challenges in handling degraded scenes, such as repetitive structures and textureless environments. Moreover, most existing methods [28, 4, 5, 27, 6] only apply to high overlap scenarios. Despite the existence of some research [1, 2] on low overlap cases, their generalization to unseen scenes is limited. Inspired by the utilization of visual fiducial markers in camera-based applications [12, 14, 13, 23] to handle degraded scenes and low overlap cases, we propose exploiting LiDAR fiducial markers to enhance multiview point cloud registration.
III Methodology
III-A Adaptive Threshold LiDAR Fiducial Marker Detection
Preliminaries. Given a 3D point cloud, IFM [19] generates a 2D intensity image by mapping each 3D point to a pixel on an image plane:
@@ III-A Adaptive Threshold LiDAR Fiducial Marker Detection
where $[u,v]^{T}$ denote the image coordinates of the projected pixel. $[\theta,\phi,r]^{T}$ denote the spherical coordinates of the 3D point. $\alpha_{a}$ and $\alpha_{i}$ represent the angular resolutions in $u$ (azimuth) and $v$ (inclination) directions, respectively. $u_{o}$ and $v_{o}$ are the offsets. The pixel value $\mathbf{I}(u,v)$ is decided by the corresponding intensity value through the color map. The ranging information is also saved for each pixel. In the real world, the intensity values of the same surface of an object could vary, leading to noise in the rendered image. To denoise the intensity image, each pixel is binarized with a threshold $\lambda$. In [19], $\lambda$ is set as a constant. Then, the embedded 2D marker detector [10, 11] is applied to localize the fiducials on the binarized image. Using the ranging information, the detected 2D fiducials are projected back into 3D space via the reverse of Eq. (1) and become 3D fiducials.
@@ III-B Problem Formulation
$\mathbf{T}^{j,*}_{i}=\underset{\mathbf{T}^{j}_{i}}{\arg\min}\sum_{s=1}^{4}\left\|\mathbf{T}^{j}_{i}\ \cdot\ ^{j}\mathbf{p}^{j,s}-_{i}\mathbf{p}^{j,s}\right\|^{2}.$ (2)
The Singular Value Decomposition (SVD) method [29] is employed to compute $\mathbf{T}^{j,*}_{i}$. We denote the set of measurements, including (1) the corner positions w.r.t. $\{M\}_{j}$, (2) the corner positions w.r.t. $\{F\}_{i}$, and (3) the marker poses w.r.t. $\{F\}_{i}$, as $\mathcal{Z}$. To register the multiview point clouds, we need to find a globally consistent pose for every scan so that the point clouds can be transformed into a complete point cloud in the global coordinate system $\{G\}$. To achieve this objective, we consider the following variables in this work: (1) the poses of point clouds w.r.t. $\{G\}$, (2) the poses of markers w.r.t. $\{G\}$, and (3) the marker corner positions w.r.t. $\{G\}$. We specify the set of variables as $\Theta$. Finally, a maximum a-posteriori (MAP) inference problem is formulated: given the measurements, $\mathcal{Z}$, the goal is to find the optimal variable set $\Theta^{*}$ that maximizes the posterior probability $P(\Theta\mid\mathcal{Z})$:
@@
III-C First-level Graph

### v3

@@ Abstract
Point cloud registration is a prerequisite for many applications in computer vision and robotics. Most existing methods focus on the registration of point clouds with high overlap. While some learning-based methods address low overlap cases, they struggle in out-of-distribution scenarios with extremely low overlap ratios. This paper introduces a novel framework dubbed L-PR, designed to register unordered low overlap multiview point clouds leveraging LiDAR fiducial markers. We refer to them as LiDAR fiducial markers, but they are the same as the popular AprilTag and ArUco markers—thin sheets of paper that do not affect the 3D geometry of the environment. We first propose an improved adaptive threshold marker detection method to provide robust detection results when the viewpoints among point clouds change dramatically. Then, we formulate the unordered multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop a framework consisting of two levels of graphs to address it. The first-level graph, constructed as a weighted graph, is designed to efficiently and optimally infer initial values of scan poses from the unordered set. The second-level graph is constructed as a factor graph. By globally optimizing the variables on the graph, including scan poses, marker poses, and marker corner positions, we tackle the MAP problem. We perform both qualitative and quantitative experiments to demonstrate that the proposed method outperforms previous state-of-the-art (SOTA) methods in addressing challenging low overlap cases. Specifically, the proposed method can serve as a convenient, efficient, and low-cost tool for applications such as 3D asset collection from sparse scans, training data collection in unseen scenes, reconstruction of degraded scenes, and merging large-scale low overlap 3D maps, which existing methods struggle with. We also collect a new dataset named Livox-3DMatch using L-PR and incorporate it into the training of the SOTA learning-based methods, which brings evident improvements for them across various benchmarks. We release the open-source code implementation and datasets at https://github.com/yorklyb/L-PR.
@@ I Introduction
Point cloud registration is a fundamental need in computer vision [1, 2] and robotics [3, 4, 5]. Most existing point registration methods [3, 6, 7, 8] focus on pairwise registration of two point clouds with high overlap, which are not robust in practical applications [2]. Some recent learning-based research proposes multiview registration methods [1, 9] and studies low overlap cases [2, 1, 9]. Despite their promising performance on benchmarks, the generalization of learning-based methods to unseen scenarios (i.e., out-of-distribution cases of training data) remains problematic. Moreover, both geometry-based methods [3, 4, 5, 10] and learning-based methods [1, 2] rely on shared geometric features between point clouds to align them. Hence, these methods struggle in scenes with scarce shared features, such as extremely low overlap cases and degraded scenes [11]. Consequently, the existing methods [9, 1, 7, 8, 3] are unsuitable for tasks such as capturing the complete 3D shape of a novel object from a sparse set of scans taken from dramatically different viewpoints, gathering data from unseen scenarios for training, reconstructing degraded scenes, or merging large-scale 3D maps with low overlap.
Reviewing the domain of 2D images, it is found that fiducial markers, such as AprilTag [12] and ArUco [13], have been widely used to tackle challenging situations, such as dramatic viewpoint changes (extremely low overlap) between frames and textureless environments [14, 15, 16]. The visual fiducial markers have rich patterns that are robust for detection and matching, thanks to elaborately designed encoding-decoding algorithms [12, 13]. They pose as thin sheets of paper that can be attached to surfaces without disrupting the 3D environment. In contrast, fiducial objects developed for LiDARs, such as calibration boards [17, 18, 19], lack encoding-decoding pattern design and often take the form of large boards mounted on tripods, thus affecting the 3D environment. LiDARTag [20] is the first work to develop a fiducial marker system akin to AprilTag [12] for LiDARs. While its patterns are compatible with AprilTag, the marker itself is still an additional object placed on a tripod. Recently, the intensity image-based LiDAR fiducial marker system (IFM) [21] proposes to utilize intensity and range images to detect markers. Thus, IFM is compatible with popular visual fiducial markers [12, 13] and inherits the virtue of not affecting the 3D environment. However, the marker detection of IFM is not robust when the viewpoints change in a wide range [21], which hinders its application to in-the-wild multiview point cloud registration. Although the development of LiDAR fiducial markers [20, 21] has advanced, their applications are limited to calibration [22] and AR [21]. The utilization of LiDAR fiducial markers in multiview point cloud registration remains an open problem.
@@
Figure 1: (a): Overview of L-PR. Given ① an unordered set of unaligned, low overlap 3D point clouds, we aim to register them into ⑤ a complete point cloud utilizing LiDAR fiducial markers (thin sheets of paper/board attached to other planes). This is achieved through the proposed adaptive threshold marker detection method and the ③④ two-level graphs. The first-level graph handles the unordered set of point clouds and calculates initial values. The second-level graph registers the point clouds by finding the optimal solution to a maximum a-posteriori problem. (b): The proposed method can serve as a convenient, efficient, and low-cost tool for diverse applications, including 3D asset collection from sparse observations, collecting training data in novel scenes, reconstruction of degraded scenes, localization in GPS-denied environments, and merging large-scale low overlap 3D maps.
In this paper, we develop a novel framework, dubbed L-PR (see Fig. 1(a)), to exploit LiDAR fiducial markers for unordered low overlap multiview point cloud registration. As seen, the LiDAR fiducial markers are thin sheets of board/paper attached to other surfaces without impacting the environment. Specifically, the LiDAR fiducial markers can be printed on paper using standard printers, offering a low-cost, easy-to-use, and reliable alternative compared to previous external sensors. Given an unordered set of low overlap point clouds, the proposed L-PR efficiently registers them in a global frame. To address the unstable marker detection of IFM [21] in the wild, we first develop an adaptive threshold marker detection method that is robust to viewpoint switching. Then, we formulate the multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop two-level graphs to tackle it. The first-level graph, constructed as a weighted graph, is designed to efficiently process the unordered set of point clouds and calculate the initial values of poses among point clouds. In particular, the weights represent the pose estimation error of each marker observation, and we obtain the optimal solutions of initial values by finding the shortest path from an anchor scan to each non-anchor scan on the graph. Given the initial values, we design the second-level graph, a factor graph, to resolve the MAP problem by globally optimizing the poses of point clouds, markers, and positions of marker corners. We conduct both qualitative and quantitative experiments to demonstrate the superiority of the proposed method over competitors [9, 1, 7, 8, 3, 16, 4]. Especially, the proposed L-PR is robust to any unseen scenarios with extremely low overlap, making it a convenient, efficient, and low-cost tool for diverse applications that pose significant challenges to existing methods. As shown in Fig. 1(b), these applications include 3D asset [25, 26, 27] collection from sparse observations, training data collection for learning-based models [9, 1] in unseen scenes, reconstruction of degraded scenes, localization in GPS-denied environments, and merging large-scale low overlap 3D maps.
We release the code implementation, datasets, and trained model weights at https://github.com/yorklyb/L-PR.
@@
  • We design a novel framework, named L-PR, for unordered low overlap multiview point cloud registration leveraging LiDAR fiducial markers. Due to the robustness of the proposed L-PR to novel scenes with extremely low overlap, L-PR can serve diverse applications that existing methods struggle with.
  • We collect a new training dataset called Livox-3DMatch using the proposed L-PR. Livox-3DMatch augments the original 3DMatch training data [6] from 14,400 pairs to 17,700 pairs (a 22.91% increase). By training on this augmented dataset, the performance of the SOTA learning-based methods [9, 1] is boosted across various benchmarks. SGHR [1] is improved by 2.90% on 3DMatch [6], 4.29% on ETH [23], and 22.72% (translation) / 11.19% (rotation) on ScanNet [24]. MDGD [9] is improved by 1.71% on 3DMatch [6], 2.89% on ETH [23], and 22.45% (translation) / 7.80% (rotation) on ScanNet [24].
@@ II Related Works
II-A LiDAR Fiducial Markers
Given the success of Structure from Motion [14, 16] and visual Simultaneous Localization and Mapping (SLAM) [15, 28, 29] approaches based on visual fiducial markers [12, 13], why do we still need to exploit LiDAR and LiDAR fiducial markers? There are three reasons: (1) Unlike image-based fiducial marker detection, which is sensitive to ambient light, LiDAR fiducial marker detection is robust to unideal illumination conditions [20, 21], such as purely dark or overexposed environments. (2) Unlike image-based fiducial marker detection [30, 31, 32], which suffers from rotational ambiguity [14, 33, 16, 15] and requires abandoning ambiguous measurements, LiDAR fiducial marker pose estimation is not affected by rotational ambiguity [20, 21] and each marker observation can be utilized. (3) Data captured by LiDARs is irreplaceable by data acquired from cameras due to the unique modality of LiDAR point cloud [34].
@@ II-A LiDAR Fiducial Markers
II-B Point Cloud Registration
Geometry-based point cloud registration methods, such as variants of iterative closest point methods [4, 5], RANSAC [10], and Teaser++ [3], mainly focus on designing efficient and robust algorithms for describing geometry and extracting geometric features (e.g., points, lines, and surfaces/normals) to find correspondence between point clouds and align them. In learning-based point cloud registration methods, neural networks are designed to learn features representing the geometry [9, 1, 35, 36, 37, 38, 6, 7, 8], and then learn to utilize these features for registering the point clouds. In addition, compared to pairwise registration methods [7, 8, 3] that only learn to align two point clouds, multiview point cloud registration methods, such as MDGD [9] and SGHR [1], also learn to refine the pose graph constructed with the pairwise registration results. While these methods demonstrate remarkable performance on benchmarks [6, 23, 24], they still face challenges in handling degraded scenes, such as repetitive structures and textureless environments. Moreover, most existing methods [39, 4, 5, 37, 6] only apply to high overlap scenarios. Despite the existence of some research [1, 2] on low overlap cases, their generalization to unseen scenes is limited. Inspired by the utilization of visual fiducial markers in camera-based applications [14, 16, 15, 28] to handle degraded scenes and low overlap cases, we propose exploiting LiDAR fiducial markers to enhance multiview point cloud registration.
III Preliminaries
III-A Three-dimensional Transformation
Suppose that $\mathbf{p}_{a}=[x,y,z]^{T}$ is a 3D point expressed in the coordinate system $\{a\}$. To express the point in another coordinate system $\{b\}$ as $\mathbf{p}_{b}=[x^{\prime},y^{\prime},z^{\prime}]^{T}$, translation and rotation transformations need be applied to $\mathbf{p}_{a}$. In particular, the extrinsic matrix, $\mathbf{T}\in\mathbb{R}^{4\times 4}$, is employed to describe the 6-DOF pose:
$\mathbf{T}=\left[\begin{array}[]{cc}\mathbf{R}&\mathbf{t}\\ \mathbf{0}^{1\times 3}&1\end{array}\right],$ (1)
where $\mathbf{t}\in\mathbb{R}^{3\times 1}$ denotes the translation vector and $\mathbf{R}\in\mathbb{R}^{3\times 3}$ denotes the rotation matrix. Afterwards, $\mathbf{p}_{b}$ is obtained as follows:
$\left[\begin{array}[]{c}\mathbf{p}_{b}\\ 1\end{array}\right]=\mathbf{T}\left[\begin{array}[]{c}\mathbf{p}_{a}\\ 1\end{array}\right].$ (2)
The operator ($\cdot$) is introduced to simplify Eq. (2) as:
$\mathbf{p}_{b}=\mathbf{T}\cdot\mathbf{p}_{a}.$ (3)
III-B Lie Group and Lie Algebra
The rotation matrix $\mathbf{R}\in SO(3)$, which is the special orthogonal group representing rotations [40]:
$SO(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}|\ \mathbf{R}\mathbf{R}^{T}=\mathbf{1},det(\mathbf{R})=1\}.$ (4)
The pose matrix $\mathbf{T}\in SE(3)$, which is the special Euclidean group representing poses [40]:
$SE(3)=\{\mathbf{T}=\left[\begin{array}[]{cc}\mathbf{R}&\mathbf{t}\\ 0&1\end{array}\right]\in\mathbb{R}^{4\times 4}|\ \mathbf{R}\in SO(3),\mathbf{t}\in\mathbb{R}^{3\times 1}\}.$ (5)
$SO(3)$ and $SE(3)$ are two specific Lie groups. Every matrix Lie group is associated with a Lie algebra [40]. The Lie algebra $\mathfrak{so}(\mathrm{3})$, associated with $SO(3)$, is defined as follows:
$SO(3)\rightarrow\mathfrak{so(\mathrm{3})}:$ (6)
$\log(\mathbf{R})=\frac{\theta}{2\mathrm{sin}(\theta)}(\mathbf{R}-\mathbf{R}^{T}),$
$\xi=\log(\mathbf{R})_{\vee},$
where $\theta=\arccos\frac{1}{2}(Trace(\mathbf{R})-1)$. $\log(\cdot)$ is the matrix logarithm and $\xi\in\mathbb{R}^{3\times 1}$ is the Lie algebra coordinates. $\vee$ is the vee map operator that finds the unique vector $\xi\in\mathbb{R}^{3\times 1}$ corresponding to a given skew-symmetric matrix $\log(\mathbf{R})\in\mathbb{R}^{3\times 3}$ [40, 28].
IV Methodology
IV-A Adaptive Threshold LiDAR Fiducial Marker Detection
Vanilla IFM. Given a 3D point cloud, the vanilla IFM [21] generates a 2D intensity image by mapping each 3D point to a pixel on an image plane:
@@ IV-A Adaptive Threshold LiDAR Fiducial Marker Detection
where $[u,v]^{T}$ denote the image coordinates of the projected pixel. $[\theta,\phi,r]^{T}$ denote the spherical coordinates of the 3D point. $\alpha_{a}$ and $\alpha_{i}$ represent the angular resolutions in $u$ (azimuth) and $v$ (inclination) directions, respectively. $u_{o}$ and $v_{o}$ are the offsets. The pixel value $\mathbf{I}(u,v)$ is decided by the corresponding intensity value through the color map. The ranging information is also saved for each pixel.
In the real world, the intensity values of the same surface of an object could vary, leading to noise in the rendered image. To denoise the intensity image, each pixel is binarized with a threshold $\lambda$. In [21], $\lambda$ is set as a constant. Then, the embedded 2D marker detector [12, 13] is applied to localize the fiducials on the binarized image. Using the ranging information, the detected 2D fiducials are projected back into 3D space via the reverse of Eq. (7) and become 3D fiducials.
@@ IV-B Problem Formulation
$\mathbf{T}^{j,*}_{i}=\underset{\mathbf{T}^{j}_{i}}{\arg\min}\sum_{s=1}^{4}\left\|\mathbf{T}^{j}_{i}\ \cdot\ ^{j}\mathbf{p}^{j,s}-_{i}\mathbf{p}^{j,s}\right\|^{2},$ (8)
where the operator ($\cdot$) is introduced in Section III-A. The Singular Value Decomposition (SVD) method [40] is employed to compute $\mathbf{T}^{j,*}_{i}$. We denote the set of measurements, including (1) the corner positions w.r.t. $\{M\}_{j}$, (2) the corner positions w.r.t. $\{F\}_{i}$, and (3) the marker poses w.r.t. $\{F\}_{i}$, as $\mathcal{Z}$. To register the multiview point clouds, we need to find a globally consistent pose for every scan so that the point clouds can be transformed into a complete point cloud in the global coordinate system $\{G\}$. To achieve this objective, we consider the following variables in this work: (1) the poses of point clouds w.r.t. $\{G\}$, (2) the poses of markers w.r.t. $\{G\}$, and (3) the marker corner positions w.r.t. $\{G\}$. We specify the set of variables as $\Theta$. Finally, a maximum a-posteriori (MAP) inference problem is formulated: given the measurements, $\mathcal{Z}$, the goal is to find the optimal variable set $\Theta^{*}$ that maximizes the posterior probability $P(\Theta\mid\mathcal{Z})$:
@@
IV-C First-level Graph